OpenStudy (anonymous):

arctan(1-x)+arctan(1+x)=arctan(1/8), solve

OpenStudy (anonymous):

apply tan on both sides and apply tan(A+B) formula u apply tan(arctan(A))= A rule to get the answer!!!!!!!

OpenStudy (anonymous):

\[\tan^{-1} (1-x)+\tan^{-1} (1+x)=\tan^{-1} (1/8)\] is that it?? but i don't know what to do next

OpenStudy (anonymous):

i got it!!! hahahahaha.. im so happy.. thanks for the help!! the answer is 4 :))

OpenStudy (anonymous):

haha... @erosha welcome :))))))))

OpenStudy (anonymous):

is my answer correct? can u check it?? :))

myininaya (myininaya):

My first step would be to subtract either arctan(1+x) or arctan(1-x Then Take tan( ) of both sides The use that one identity: \[\tan(a \pm b)=\frac{\tan(a) \pm \tan(b)}{1 \mp \tan(a)\tan(b)}\]

myininaya (myininaya):

Then* not the

myininaya (myininaya):

The rest is algebra

myininaya (myininaya):

Because as @sriramkumar said tan(arctan(a))=a

OpenStudy (anonymous):

1-x+1+x/(1-(1-x)(1+x))= 1/8 2/1-(1-x^2)= 1/8 2/x^2 = 1/8 x^2= 8*2=16 x=+/-4 yep its correct!!!!!! @eroshea